# Deep Reasoning Translation via Reinforcement Learning

### v1

@@ Abstract
Recently, deep reasoning LLMs (e.g., OpenAI o1/o3 and DeepSeek-R1) have shown promising performance in various complex tasks. Free translation is an important and interesting task in the multilingual world, which requires going beyond word-for-word translation and taking cultural differences into account. This task is still under-explored in deep reasoning LLMs. In this paper, we introduce DeepTrans, a deep reasoning translation model that learns free translation via reinforcement learning. Specifically, we carefully build a reward model with pre-defined scoring criteria on both the translation results and the thought process. Given the source sentences, the reward model teaches the deep translation model how to think and free-translate them during reinforcement learning. In this way, training DeepTrans does not need any labeled translations, avoiding the human-intensive annotation or resource-intensive data synthesis. Experimental results show the effectiveness of DeepTrans. Using Qwen2.5-7B as the backbone, DeepTrans improves performance by 16.3% in literature translation, and outperforms strong deep reasoning baselines as well as baselines that are fine-tuned with synthesized data. Moreover, we summarize the failures and interesting findings during our RL exploration. We hope this work could inspire other researchers in free translation. [^1] https://github.com/krystalan/DRT
Deep Reasoning Translation via Reinforcement Learning
Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,withtomzhou}@tencent.com
@@ 1 Introduction
In recent years, deep reasoning LLMs OpenAI (2024b); Guo et al. (2025) have shown remarkable performance in tasks ranging from math Chen et al. (2025b); Li et al. (2025), codes Zhang et al. (2024), question-taking Guan et al. (2025), etc.
@@
Some researchers bring the success of deep reasoning LLMs to neural machine translation (MT). Marco-o1 Zhao et al. (2024) is introduced for open-ended text generation, and it uses illustrative examples to show the effectiveness of long chain-of-thought (CoT) in colloquial and slang MT. Wang et al. (2024) further systematically study the long CoT in literature MT. They find the literature sentences with metaphors or similes generally require cultural background to understand. Based on this insight, Wang et al. (2024) propose DRT(-o1) LLMs to translate the literature text from English to Chinese with step-by-step reasoning. The DRT LLMs are trained from synthesized long CoT translation data. Chen et al. (2025a) conduct empirical studies on how deep reasoning LLMs work on MT. They verify the strengths of deep reasoning LLMs in historical and cultural translation, but also point out their rambling issues, i.e., LLMs do not follow the MT instruction and fail to translate. Liu et al. (2025) also briefly show the cases of deep reasoning LLMs in MT. More recently, R1-T1 He et al. (2025) is presented, which is the first attempt to employ reinforcement learning (RL) in deep reasoning MT models. Specifically, it uses COMET Rei et al. (2020) score as the reward signal to optimize MT LLMs via modified REINFORCE++ (a RL training algorithm) Hu (2025).
@@
Meanwhile, RL has been demonstrated to have a strong ability to create deep reasoning LLMs Guo et al. (2025); Yu et al. (2025). LLMs can be equipped with powerful reasoning ability solely based on RL training, and such an ability could reveal promising performance on complex tasks (such as math and coding problems). However, RL is still under-explored in deep reasoning MT LLMs. Marco-o1 and DRT are solely trained via supervised fine-tuning (SFT), which requires high computation resources to synthesize data. Though R1-T1 leverages RL during its creation, using the COMET score as the reward signals is not well calibrated. As pointed by Liu et al. (2025), DeepSeek-R1 (a deep reasoning LLM) achieves better translation but a lower COMET compared with DeepSeek-V3 (a non-reasoning LLM). Besides, in some domains, COMET loses its effectiveness in MT evaluation and shows a poor correlation with humans Karpinska and Iyyer (2023).
@@
In this paper, our research goal is to explore how to improve the free translation ability of deep reasoning LLMs via reinforcement learning. Given the strong ability of GRPO Shao et al. (2024), we adopt it as the RL training framework. Consequently, the core problem is how to employ an effective reward model during RL training. There might be three mainstream types of reward models in previous LLM work: i) using widely-adopted MT metrics such as BLEU Papineni et al. (2002), BLEURT Sellam et al. (2020), COMET Rei et al. (2020) and CometKiwi Rei et al. (2022); ii) training a reward model with preference data Stiennon et al. (2020); Nakano et al. (2021), and use the reward model to inference a scalar reward during RL training; iii) designing rule-based rewards Guo et al. (2025). However, these reward models might lose their effectiveness or require high-quality annotated data in MT, hindering their usage. In detail, for i), the reference-based BLEU, BLEURT and COMET are unsuitable for literature MT since the high-quality translation data is extremely difficult to collect. Though CometKiwi is a reference-free metric, its effectiveness is limited in the literature domain, achieving low correlation with humans Wang et al. (2024). For ii), training a reward model requires large-scale annotated preference data, which is also difficult to collect due to the expensive annotation costs. For iii), the rule-based rewards are suitable for tasks whose answers are easy to verify, such as math and code problems Swamy et al. (2025). In MT, we cannot design simple rules to judge the quality of translations.
@@
In view of the strong ability of LLM-as-a-judge Li et al. (2024); Wang et al. (2023); Kocmi and Federmann (2023), we decide to use an advanced instruct LLM as the reward model. Specifically, we carefully design pre-defined scoring criteria on both the translation results and the thought process. For the generation results of deep reasoning MT LLMs, the reward model takes the pre-defined criteria into account to provide a discrete reward (e.g., 3-point or 100-point scores) for RL training. In this manner, we avoid collecting human-annotated preference data for training the reward model. In addition, the effectiveness of the reward model could also be ensured when we adopt the state-of-the-art LLM. We conduct extensive experiments on literature MT, using the above RL strategy to train our DeepTrans model (with the backbone of Qwen2.5-7B). Experimental results show the effectiveness of our method, DeepTrans improves performance by 16.3% and outperforms strong deep reasoning baselines (e.g., QwQ-32B-preview and Marco-o1). Besides, only training with source sentences, DeepTrans outperforms DRT-7B (which is trained on synthesized long thought and translation data) in terms of both automatic metrics and GPT-4o evaluation. Moreover, we share the failures in our pilot RL experiments, and summarize the critical component in reward modeling. We hope our experiments could inspire other researchers in free translation.
Our main contributions are concluded as follows:
@@
  • We propose DeepTrans, which aims at enhancing the free translation ability of deep reasoning LLMs via reinforcement learning. To this end, we use an advanced LLM as the reward model, and carefully design scoring criteria on both translations and thought process.
@@
  • Experimental results on literature translation verify the effectiveness of our DeepTrans. Only training with the source sentences in the RL framework, DeepTrans outperforms strong deep reasoning baselines and LLMs that are fine-tuned with synthesized literature MT data.
@@
Figure 1: The overview of DeepTrans RL training.
@@ 2 DeepTrans
In this section, we introduce DeepTrans. As illustrated in Figure 1, there are three rewards during RL training: format reward, thought reward and translation reward. The first one is obtained via both rules and LLMs, while the latter two rewards are judged by LLMs. We first discuss the rewards designed in the RL framework (§ 2.1) and then provide the training details of DeepTrans (§ 2.2).
@@ 2.1 Reward Modeling
Given a source sentence, DeepTrans first thinks and analyzes how to translate the sentence, and finally provides the translation result. We design the format of model generation as “<think> [thought] </think> [translation]”, where “<think>” and “</think>” are two special tokens to indicate the boundary of thought content. “[thought]” and “[translation]” denote the content of thought and translation, respectively. Based on the model generation, we design the following rewards:
Format Reward: We use a regular expression to judge whether the generation format is correct. Besides, in pilot experiments, we find that there might be some explanations in the translation results (i.e., [translation]). To avoid it, we employ DeepSeek-v3 Liu et al. (2024) to judge whether the translation results only contain translations. The judgment prompt is shown as follows:
@@
If both (1) the generation format is correct, determined by the regular expression, and (2) the translation result only contains translation determined by DeepSeek-v3, we regard the format as correct; otherwise, it is incorrect:
$r_{\text{format}}=\begin{cases}1&\text{if format is correct}\\ 0&\text{if format is incorrect}\end{cases}$ (1)
Thought Reward: The thought process is important to guide the final translation. To reward logical thought processes, we adopt DeepSeek-v3 to provide the thought reward with a 3-point scale.
A translation question requires translating a given text from [src lang] into [trg lang]. The given text is as follows: <text> {src} </text> Someone did this translation question, and began to think how to translate: <think> {think} </think> Please judge whether there is a detailed analysis of the given text in this thinking process: 1. No analysis: Only very shallow thinking was done, and no detailed analysis of the given text was carried out. 2. Slight analysis: The given text was analyzed in detail, and how to translate it was discussed in detail. 3. Detailed analysis: The given text was analyzed in detail, and various translation possibilities were discussed in detail, and trade-offs were made. Please directly output your judgment results, such as: “no analysis”, “slight analysis” or “detailed analysis”
where “{think}” denotes the thought content. The thought reward can be further defined as follows:
$r_{\text{thought}}=\begin{cases}2&\text{if}\operatorname{DS-v3}^{\text{th}}(\text{src},\text{think})=\text{detailed analysis}\\ 1&\text{if}\operatorname{DS-v3}^{\text{th}}(\text{src},\text{think})=\text{slight analysis}\\ 0&\text{if}\operatorname{DS-v3}^{\text{th}}(\text{src},\text{think})=\text{no analysis}\\ \end{cases}$ (2)
where $\operatorname{DS-v3}^{\text{th}}$ denotes using DeepSeek-v3 as the thought reward scorer.
@@
SYSTEM PROMPT: Please evaluate the following Chinese translation of an English text. Rate the translation on a scale of 0 to 100, where: - 10 points: Poor translation; the text is somewhat understandable but contains significant errors and awkward phrasing that greatly hinder comprehension for a Chinese reader. - 30 points: Fair translation; the text conveys the basic meaning but lacks fluency and contains several awkward phrases or inaccuracies, making it challenging for a Chinese reader to fully grasp the intended message. - 50 points: Good translation; the text is mostly fluent and conveys the original meaning well, but may have minor awkwardness or slight inaccuracies that could confuse a Chinese reader. - 70 points: Very good translation; the text is smooth and natural, effectively conveying the intended meaning, but may still have minor issues that could slightly affect understanding for a Chinese reader. - 90 points: Excellent translation; the text is fluent and natural, conveying the original meaning clearly and effectively, with no significant issues that would hinder understanding for a Chinese reader. Please provide the reason first, followed by a score. Format your evaluation in the JSON structure below: {"reason": "reason for the score", "score": int} USER PROMPT: <text> {src} </text> <translation> {trans} </translation>
In this manner, DeepSeek-v3 will generate its judgment for a translation and provide its reward with a 100-point scale:
$r_{\text{trans}}=\operatorname{DS-v3}^{\text{tr}}(\text{src},\text{trans})$ (3)
where $\operatorname{DS-v3}^{\text{tr}}$ denotes using DeepSeek-v3 as the translation reward scorer.
@@
$r_{\text{all}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ r_{\text{trans}}+\alpha\times r_{\text{thought}}&\text{if }r_{\text{format}}\neq 0\end{cases}$ (4)
where $\alpha$ is a trade-off hyperparameter between $r_{\text{trans}}$ and $r_{\text{thought}}$ ($\alpha$ is set to 20 in experiments).
@@ 2.2 Training Details
Cold Start SFT. We use a non-reasoning LLM, i.e., Qwen2.5-7B-Instruct Yang et al. (2024), as the backbone of DeepTrans. To adapt DeepTrans to the deep reasoning MT task, we first use DeepSeek-R1 Guo et al. (2025) to generate seed translation samples in the general domain (instead of the literature domain). The seed translation samples follow our designed format, and are further used to few-shot SFT the backbone model (named cold start SFT). It is also worth noting that the goal of cold start SFT is not to teach the model free translation but the general translation with long thought.
RL Training. In view of the strong ability of GRPO Shao et al. (2024), which discards the critic model in PPO Schulman et al. (2017). For the policy model $\pi$, given it a source sentence $s$, GRPO first samples a number of generations $\{g_{1},g_{2},...,g_{n}\}$ based on $\pi$, where each $g_{i}$ involves a thought process and the corresponding translation result. Then, GRPO optimizes the policy model $\pi^{\prime}$ by maximizing the following objective:
$\frac{1}{n}\sum_{1}^{n}(\operatorname{min}(\nabla_{\pi}A_{i},\operatorname{clip}(\nabla_{\pi},1-\epsilon,1+\epsilon)A_{i})-\beta\mathcal{D}$ (5)
$\nabla_{\pi}=\frac{\pi^{\prime}(g_{i}|s)}{\pi(g_{i}|s)}$ (6)
$\mathcal{D}=\mathbb{D}_{kl}(\pi^{\prime}||\pi_{ref}))$ (7)
@@
$A_{i}=\frac{r_{\text{all}}^{i}-{\operatorname{mean}(\{r_{\text{all}}^{1},r_{\text{all}}^{2},\cdots,r_{\text{all}}^{n}\})}}{{\operatorname{std}(\{r_{\text{all}}^{1},r_{\text{all}}^{2},\cdots,r_{\text{all}}^{n}\})}}.$ (8)

### v2

# DeepTrans: Deep Reasoning Translation via Reinforcement Learning

@@ Abstract
Recently, deep reasoning LLMs (e.g., OpenAI o1 and DeepSeek-R1) have shown promising performance in various downstream tasks. Free translation is an important and interesting task in the multilingual world, which requires going beyond word-for-word translation. However, the task is still under-explored in deep reasoning LLMs. In this paper, we introduce DeepTrans, a deep reasoning translation model that learns free translation via reinforcement learning (RL). Specifically, we carefully build a reward model with pre-defined scoring criteria on both the translation results and the thought processes. The reward model teaches DeepTrans how to think and free-translate the given sentences during RL. Besides, our RL training does not need any labeled translations, avoiding the human-intensive annotation or resource-intensive data synthesis. Experimental results show the effectiveness of DeepTrans. Using Qwen2.5-7B as the backbone, DeepTrans improves performance by 16.3% in literature translation, and outperforms strong deep reasoning LLMs. Moreover, we summarize the failures and interesting findings during our RL exploration. We hope this work could inspire other researchers in free translation. [^1] https://github.com/krystalan/DRT
@@ 1 Introduction
Recently, deep reasoning LLMs OpenAI (2024b); Guo et al. (2025) have shown remarkable performance in tasks ranging from math Chen et al. (2025b); Li et al. (2025), to coding Zhang et al. (2024), question-taking Guan et al. (2025), etc.
@@
Some researchers bring the success of deep reasoning LLMs to neural machine translation (MT), with a particular focus on free translation rather than word-for-word translation Zhao et al. (2024); Wang et al. (2025); Liu et al. (2025). Free translation is a translation approach that allows more flexibility to adapt the text to the target language, taking into account cultural nuances and making the text more natural and understandable for the target audience Barbe (1996); Chen et al. (2018). Given that, free translation is well-suited for deep reasoning LLMs to perform. Marco-o1 Zhao et al. (2024) aims to extend deep reasoning LLMs into general domains where clear standards may be lacking. It uses illustrative examples to show the effectiveness of long chain-of-thought (CoT) in colloquial and slang MT. Wang et al. (2025) further systematically study the long CoT in literature MT. They find the literature sentences with metaphors or similes generally require cultural background to understand. Based on this insight, Wang et al. (2025) propose deep reasoning translation (DRT) models to translate the literature text from English to Chinese with step-by-step reasoning. The models are trained from synthesized long CoT translation data. Chen et al. (2025a) conduct empirical studies on how deep reasoning LLMs work on MT. They verify the strengths of deep reasoning LLMs in historical and cultural translation, but also point out their issues, e.g., LLMs do not follow the instruction and fail to translate. More recently, R1-T1 He et al. (2025) is presented, which is the first attempt to employ reinforcement learning (RL) in deep reasoning LLMs. Specifically, it uses COMET Rei et al. (2020) score as the reward signal to optimize MT LLMs via modified REINFORCE++ (a RL training algorithm) Hu (2025). MT-R1 Feng et al. (2025) uses a combination of BLEU and CometKiwi as the reward signal to optimize MT LLMs, and adopts GRPO Shao et al. (2024) as the RL algorithm.
@@
Meanwhile, RL has been verified to have a strong ability in deep reasoning LLMs, and LLMs can be equipped with powerful reasoning ability solely based on RL Guo et al. (2025); Yu et al. (2025). However, RL is still under-explored in deep reasoning MT. In detail, Marco-o1 and DRT are only trained via supervised fine-tuning (SFT). Though R1-T1 and MT-R1 adopt RL training, using the COMET and CometKiwi as the rewards is not well calibrated. As pointed out by Liu et al. (2025), deep reasoning MT LLMs might achieve better translation but a lower COMET. Besides, in literature domains, COMET and CometKiwi lose their effectiveness as evaluation metrics and show a poor correlation with human judgement Karpinska and Iyyer (2023); Wang et al. (2025).
@@
In this paper, our research goal is to explore how to improve the free translation ability of deep reasoning LLMs via reinforcement learning. Following Wang et al. (2025), our explorations focus on the literature domain, where free translation is essential for addressing cultural differences. Subsequently, we focus on the reward modeling, and study how to employ an effective reward model during RL training. There are three mainstream types of reward models in previous work: i) using MT metrics such as BLEU Papineni et al. (2002), COMET Rei et al. (2020) and CometKiwi Rei et al. (2022); ii) training a reward model with preference data Stiennon et al. (2020); Nakano et al. (2021), and using the reward model to infer a scalar reward during RL training; iii) designing rule-based rewards Guo et al. (2025). However, these rewards might lose their effectiveness or require high-quality annotated data in MT, hindering their usage. In detail, for i), the reference-based BLEU and COMET are unsuitable for literature MT since the high-quality translation data is difficult to collect. Though CometKiwi is a reference-free metric, its effectiveness is limited in the literature domain Wang et al. (2025). For ii), training a reward model requires large-scale preference data, which is difficult to collect due to the expensive annotation costs. For iii), the rule-based rewards are suitable for tasks whose answers are easy to verify, e.g., math and code problems Swamy et al. (2025). In MT, we cannot design simple rules to judge the quality of translations.
@@
In view of the strong ability of LLM-as-a-judge Wang et al. (2023); Kocmi and Federmann (2023); Li et al. (2024), we decide to use an advanced LLM, i.e., DeepSeek-v3 (671B) Liu et al. (2024), as the reward model. Specifically, we carefully design pre-defined scoring criteria on both the translation results and the thought processes. For the generation of deep reasoning MT LLMs, the reward model takes the pre-defined criteria into account to provide a discrete reward (e.g., 3-point or 100-point scores). In this manner, we avoid collecting human-annotated preference data for training the reward model. In addition, the effectiveness of the reward model could also be ensured when we adopt the state-of-the-art LLM. We conduct extensive experiments on literature MT, using the above RL strategy to train DeepTrans-7B (with the backbone of Qwen2.5-7B). Experiments show the effectiveness of our method, DeepTrans-7B improves performance by 16.3% and outperforms strong deep reasoning baselines (e.g., QwQ-32B-preview). Besides, only training with source sentences, DeepTrans-7B outperforms DRT-7B (which is trained on synthesized long thought and translation data) in terms of both automatic metrics and GPT-4o evaluation. Moreover, we share the failures in our pilot RL experiments and summarize the critical component in reward modeling.
Our contributions are as follows:
@@
  • We propose DeepTrans, which aims to enhance the free translation ability of deep reasoning LLMs via RL. In detail, we use an LLM as the reward model, and carefully design scoring criteria on both translations and thought processes.
@@
  • Experimental results verify the effectiveness of our DeepTrans. Only training with the source sentences, DeepTrans outperforms strong deep reasoning baselines and LLMs that are fine-tuned with synthesized MT data.
@@
2 Related Work
Deep Reasoning LLMs. In recent years, deep reasoning LLMs have pioneered a growing research in long chain-of-thought (CoT) reasoning. Different from the non-reasoning-oriented LLMs, deep reasoning LLMs involve a more detailed, iterative process of exploration and reflection within a given problem space Chen et al. (2025b); Li (2025). Many studies investigate the mathematical reasoning, programming tasks, and multidisciplinary knowledge reasoning capabilities of deep reasoning LLMs, and achieve promising performance Yu et al. (2024); Sun et al. (2025); Yax et al. (2024); Li et al. (2025); Guan et al. (2025); Jin et al. (2025). Some researchers investigate the MT capability of deep reasoning LLMs. Zhao et al. (2024) and Liu et al. (2025) briefly show that the long CoT reasoning helps the model to reach more idiomatic translations. DRT Wang et al. (2025) is further proposed in literary translation, and it is trained with synthesized SFT data. More recently, R1-T1 He et al. (2025) and MT-R1 Feng et al. (2025) employ RL to improve the translation ability of deep reasoning LLMs. However, R1-T1 and MT-R1 use the traditional MT metrics, i.e., BLEU Papineni et al. (2002), COMET Rei et al. (2020) and COMETKiwi Rei et al. (2022), as the RL reward. Different from them, we employ an advanced LLM with pre-defined scoring criteria to qualify the quality of both the translation and the thought process as the reward. In this way, the flaws of traditional metrics Karpinska and Iyyer (2023); Wang et al. (2025) can be avoided, and the strong ability of LLM-as-a-judge can be utilized to guide the RL training process effectively.
RL in Traditional MT. Machine translation via reinforcement learning has been explored before the deep reasoning LLM era. Early work tries to train MT models via optimizing BLEU scores Ranzato et al. (2016); Shen et al. (2016); Bahdanau et al. (2017). Wu et al. (2016) design GLEU scores as the rewards to deal with the drawbacks of BLEU in single sentence evaluation. Wu et al. (2018) propose a method to involve large-scale monolingual data during RL training. Choshen et al. (2020) show the challenges of optimizing MT models via RL, e.g., sparse reward signals and high-dimensional action space. Kiegeland and Kreutzer (2021) provide further analyses on these challenges. Kang et al. (2020) study document-level MT, and they propose a new method to both select context and translate sentences via RL.
@@ 3 DeepTrans
In this section, we introduce DeepTrans. As illustrated in Figure 1, there are three types of rewards: format reward, thought reward and translation reward. We first discuss the rewards designed in the RL framework (§ 3.1) and then provide the training details of DeepTrans (§ 3.2).
@@ 3.1 Reward Modeling
Given a source sentence, DeepTrans first thinks about how to translate the sentence, and then provides the translation result. We design the format of model generation as “<think> [thought] </think> [translation]”, where “<think>” and “</think>” are two special tokens to indicate the boundary of thought content. “[thought]” and “[translation]” denote the content of thought and translation, respectively. Based on the model generation, we design the following rewards:
Format Reward: We use a regular expression to judge whether the generation format is correct. Besides, in pilot experiments, we find that there might be some explanations in the translation results. To avoid it, we employ DeepSeek-v3 Liu et al. (2024) to judge whether the translation results only contain translations. The judgment prompt is shown as follows:
@@
Figure 1: The overview of DeepTrans RL training.
If both (a) the generation format is correct (determined by the regular expression), and (b) the translation result does not contain any explanations (determined by DeepSeek-v3), we regard the format as correct; otherwise, it is incorrect:
$\small r_{\text{format}}=\begin{cases}1&\text{if format is correct}\\ 0&\text{if format is incorrect}\end{cases}$ (1)
Thought Reward: The thought process is important to guide the final translation. To reward meaningful thought processes, we adopt DeepSeek-v3 to provide the thought reward with a 3-point scale:
A translation question requires translating a given text from [src lang] into [trg lang]. The given text is as follows: <text> {src} </text> Someone did this translation question, and began to think how to translate: <think> {think} </think> Please judge whether there is a detailed analysis of the given text in this thinking process: 1. A lack of analysis: Only very shallow thinking was done, and no detailed analysis of the given text was carried out. 2. Slight analysis: The given text was analyzed in detail, and how to translate it was discussed in detail. 3. Detailed analysis: The given text was analyzed in detail, and various translation possibilities were discussed in detail, and trade-offs were made. Please directly output your judgment results, such as: “a lack of analysis”, “slight analysis” or “detailed analysis”
where “{think}” denotes the thought content. Subsequently, we define the thought reward as:
[…]
where $\operatorname{v3}^{\text{th}}(\cdot,\cdot)$ denotes using DeepSeek-v3 as the thought reward scorer.
@@
SYSTEM PROMPT: Please evaluate the following Chinese translation of an English text. Rate the translation on a scale of 0 to 100, where: - 10 points: Poor translation; the text is somewhat understandable but contains significant errors and awkward phrasing that greatly hinder comprehension for a Chinese reader. - 30 points: Fair translation; the text conveys the basic meaning but lacks fluency and contains several awkward phrases or inaccuracies, making it challenging for a Chinese reader to fully grasp the intended message. - 50 points: Good translation; the text is mostly fluent and conveys the original meaning well, but may have minor awkwardness or slight inaccuracies that could confuse a Chinese reader. - 70 points: Very good translation; the text is smooth and natural, effectively conveying the intended meaning, but may still have minor issues that could slightly affect understanding for a Chinese reader. - 90 points: Excellent translation; the text is fluent and natural, conveying the original meaning clearly and effectively, with no significant issues that would hinder understanding for a Chinese reader. Please provide the reason first, followed by a score. Format your evaluation in the JSON structure below: {"reason": "reason for the score", "score": int}
USER PROMPT: <text> {src} </text> <translation> {trans} </translation>
In this way, DeepSeek-v3 will generate its judgment for the translation result and provide a reward score on a 100-point scale:
$\small r_{\text{trans}}=\operatorname{v3}^{\text{tr}}(\text{src},\text{trans})$ (3)
where $\operatorname{v3}^{\text{tr}}(\cdot,\cdot)$ denotes using DeepSeek-v3 as the translation reward scorer.
@@
$\small r_{\text{all}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ r_{\text{trans}}+\alpha\times r_{\text{thought}}&\text{if }r_{\text{format}}\neq 0\end{cases}$ (4)
where $\alpha$ is a trade-off hyperparameter between $r_{\text{trans}}$ and $r_{\text{thought}}$. Note that DeepSeek-v3 is used in $r_{\text{trans}}$ / $r_{\text{thought}}$, and its effectiveness in MT evaluation is also verified by Larionov et al. (2025).
@@ 3.2 Training Details
Cold Start SFT. We use a non-reasoning LLM, i.e., Qwen2.5-7B-Instruct Yang et al. (2024), as the backbone of DeepTrans. To adapt DeepTrans to the deep reasoning MT, we first use DeepSeek-R1 Guo et al. (2025) to generate seed translation samples in the general domain (instead of the literature domain). These samples follow our designed format, and are used to few-shot SFT the backbone model (named cold start SFT). The goal of cold start SFT is not to teach the model free translation, but the general translation with long thought.
RL Training. In view of the strong ability of GRPO Shao et al. (2024), we adopt it in our RL training. For the policy model $\pi$, given it a source sentence $s$, GRPO first samples a number of generations $\{g_{1},g_{2},...,g_{n}\}$ based on $\pi$, where each $g_{i}$ involves a thought process and the translation result. Then, GRPO optimizes the policy model $\pi^{\prime}$ by maximizing the following objective:
$\small\frac{1}{n}\sum_{1}^{n}(\operatorname{min}(\nabla_{\pi}A_{i},\operatorname{clip}(\nabla_{\pi},1-\epsilon,1+\epsilon)A_{i})-\beta\mathcal{D}$ (5)
$\small\nabla_{\pi}=\frac{\pi^{\prime}(g_{i}|s)}{\pi(g_{i}|s)}$ (6)
$\small\mathcal{D}=\mathbb{D}_{kl}(\pi^{\prime}||\pi_{ref}))$ (7)
@@
$\small A_{i}=\frac{r_{\text{all}}^{i}-{\operatorname{mean}(\{r_{\text{all}}^{1},r_{\text{all}}^{2},\cdots,r_{\text{all}}^{n}\})}}{{\operatorname{std}(\{r_{\text{all}}^{1},r_{\text{all}}^{2},\cdots,r_{\text{all}}^{n}\})}}.$ (8)